Little Dome C - Beyond EPICA Drilling Site (75.29917 degS, 122.44516 degE)

Situation Report #11, 12. December 2021

Personnel @LDC: Saverio Panichi (ENEA, Camp Leader), Carlo Barbante (ISP-CNR, PI), Olivier Alemany (IGE), Giacomo Bonanno (ENEA), Olivier Delanoe (IPEV), Romain Duphil (IPEV), Matthias Hüther (AWI), Calogero Monaco (ENEA), Anthony Pauty (IPEV), Philippe Possenti (IPEV), Michele Scalet (ENEA), Jakob Schwander (unibe), Thomas Stocker (unibe), Remo Walther (unibe)

Personnel @DC: Gregory Teste (IGE)

Weather at LDC 9 pm: sunny

Meteo at DC 9 pm: T = −31°C, Wind = 15 knt, Wind Chill T = −46°C

A sunny Sunday, but high winds did not invite for a stroll around Little Dome C. Those who worked outside suffered hard conditions. Considering the wind chill, a value that expresses the perceived temperature, the BEOIC team was exposed to −46°C today.

Giacomo and the electricians have worked on the connection of the satellite dish with the main power and attached the coax cable to the main unit in the recreation tent. The work outside was particularly taxing due to the high winds today (15 knots). Also, the antenna was placed on an elevated snow platform for improved satellite communication. The platform was custom-made by Michele diligently maneuvering the Pistenbully and its blade.

The second step of the reaming was continued today by Philippe and Carlo. Olivier A. and Mathias have installed a window into the Vissmann cabin for drilling fluid recycling and worked further on the electrical infrastructure.

BEOIC drilling depth at the end of the day: 128.0 m

RADIX firn drilling made a substantial progress owing to the stretched tube which showed much less friction in the bore hole. After several runs we have reached a depth of 50 m. Tomorrow the nylon tube will undergo another stretching under elevated temperature.

Radix drilling depth: 50 m

Thomas Stocker and Carlo Barbante